How many total atoms are in 0.320 g of P2O5?
1
Expert's answer
2019-10-04T04:39:49-0400

Molar mass of "P_2O_5=142 gm"

Number of moles of "P_2O_5 \ in \ 0.320 \ gm=\frac{0.32}{142}"

Number of atoms in "P_2O_5 \ sample=7 \times N_A\times \frac{0.32}{142}=9.5\times 10^{21}"

Where "N_A" is avagadro number("6.023 \times 10^{23})"